4. Calculate and describe how to make a 100 mL solution of 3 mM potassium hexacyanoferrate (III) (329.24 g/mol) in 0.1 M potassium chloride.
Answer #1

potassium hexacyanoferrate(iii) is soluble in KCl

mass of KCl required to prepare1L of 0.1M solution=(0.1 mol/L)*1L*(molar mass of KCl)=(0.1 mol/L)*1L*(74.551 g/mol)=7.4551 g

Dissolve 7.4551g KCl in 1L of distilled water ,and separate 100 ml of this solution.separate 100 ml of this KCl solution and add the required mass of potassium hexacyanoferrate(iii)

concentration of potassium hexacyanoferrate(iii) required=3*10^-3 mol/L

mol of potassium hexacyanoferrate(iii) in 100 ml=(3*10^-3 mol/L)*0.1L=3*10^-4 mol

mass of potassium hexacyanoferrate(iii)=3*10^-4 mol *329.24g/mol=0.09877 g

weigh 0.09877 g of potassium hexacyanoferrate(iii) and add to the above prepared KCl solution.

Thus, the solution is ready (100ml of 3mM mass of potassium hexacyanoferrate(iii) in 0.1M KCl)